A stone is thrown horizontally off a 25 m high building. It hits the ground a distance of 20 m from the foot of the building. What is the initial velocity of the stone?

1
Expert's answer
2022-04-17T17:08:10-0400

Since the velocity in horizontal direction does not change during the motion, the initial velocity can be found as follows:


"v_0 = \dfrac{d}{t}"

where "d = 20 m" is the distance travelled in horizontal direction, and "t" is the time of fall.

From the kinematic equation


"h = \dfrac{gt^2}{2}"

where "h=25m" is the distance travelled in vertical direction, one can express the time of fall:


"t = \sqrt{\dfrac{2h}{g}}"

Substituting it into the expression for velocity, obtain:


"v_0 = d\sqrt{\dfrac{g}{2h}} =20m\cdot \sqrt{\dfrac{9.8m/s^2}{2\cdot 25m}} \approx 8.9m/s"

Answer. 8.9 m/s.